124. Marriage penalty eliminated. The value of the expression

4220 + 0.25(x -30,650)
Is the 2006 federal income tax for a single taxpayer with
Taxable income of x dollars, where x is over $30,650 but
Not over $74,200.

a) Simplify the expression.
b) Find the amount of tax for a single taxpayer with
Taxable income of $40,000.
c) Who pays more, a married couple with a joint taxable
Income of $80,000 or two single taxpayers with
Taxable incomes of $40,000 each?

Answer by stanbon(75887)

The value of the expression 4220 + 0.25(x -30,650)
Is the 2006 federal income tax for a single taxpayer with
Taxable income of x dollars, where x is over $30,650 but
Not over $74,200.
a) Simplify the expression.
4220 + 0.25(x -30,650)
= 4220 + (x/4) - 7662.5
= (x/4) - 3442.50
-----------------------------
b) Find the amount of tax for a single taxpayer with
Taxable income of $40,000.
Ans: 40,000/4 - 3442.5
= 10,000-3442.5
---
= 6557.50
=================================

c) Who pays more, a married couple with a joint taxable
Income of $80,000 or two single taxpayers with
Taxable incomes of $40,000 each?
---
I'll leave that to you.
=======
Cheers,
Stan H.